OpenStudy (anonymous):

Janie flips a coin 70 times and records if it comes up heads. If getting heads is a success, what is the probability of a failure?

OpenStudy (anonymous):

Well since there are only two sides, the chance of not getting heads 70 times is 1/2 of 70. So do 70/2 or 1/2 * 70 in order to find the answer. Then you can put the answer (x) over 70 and then simplify to 1/2.

OpenStudy (anonymous):

so wats that answer

OpenStudy (anonymous):

I cannot give you the direct answer. You must work it out for yourself.

OpenStudy (anonymous):

i got 35

OpenStudy (anonymous):

In order to get the answer, you must find 1/2 of 70 - the probability will be shown as a percent.

OpenStudy (anonymous):

So if you divide by two, that means 35 is half of 70

OpenStudy (anonymous):

is it 0.2

OpenStudy (anonymous):

So the fraction is 1/2

OpenStudy (anonymous):

So as a decimal, it is 0.5

OpenStudy (mathmale):

It's possible that not all the info given in this problem statement is necessary. The chances of getting a head is simply 0.5. What are the chances of getting a tail (not getting a head)? Please note carefully: 35 is NOT a viable probability. Probabilities are always between 0 and +1. So that rules out 35 as an answer. Again: What are the chances of not getting a head?

OpenStudy (anonymous):

When creating percents we move the decimal point 2 places to the right

OpenStudy (anonymous):

So it is a 50% chance

OpenStudy (mathmale):

Chance of what? Please be specific.

OpenStudy (anonymous):

Do you understand? Because you need to learn not cheat.

OpenStudy (anonymous):

50% chance of not getting a head

OpenStudy (anonymous):

ok can you help with this question

OpenStudy (anonymous):

Thirty-three percent of all students at a high school like beef stew. Out of 300 students, 40 are selected randomly and asked whether they liked beef stew. If liking beef stew is a success, what is the probability of a failure for this binomial experiment?

OpenStudy (anonymous):

40 over 300 times 100. The answer must be subtracted from 300.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Look, sexy, it is important that you are learning and do not just want answers. Do you understand probability or no?

OpenStudy (anonymous):

no

OpenStudy (mathmale):

Paraphrase the question as follows: "The probability that a student selected at random at a certain high school likes beef stew is about 1/3 (33.3%). What is the probability that a student selected at random does NOT like beef stew?"

OpenStudy (mathmale):

Another way to put this would be: "In a situation in which there are only two possible outcomes to an experiment, the probability of success 1/3. What is the probability of failure in this experiment?"

OpenStudy (anonymous):

Okay, um, I am going to leave you in the hands of @mathmale but contact your teacher for more questions or go to Spy Guys 6 - a good math website

OpenStudy (anonymous):

i dont know

OpenStudy (mathmale):

If P(success) in a binomial experiment is 1/3, then P(failure) is simply 1-1/3 = 2/3.

OpenStudy (mathmale):

What specifically do you want to know? If you're not familiar with the terms I'm using, we could look them up on the Internet. it's difficult to respond to "what I dot get".

OpenStudy (anonymous):

how to get the answer

OpenStudy (mathmale):

In this case, all you really need to know is in what I typed earlier: "If P(success) in a binomial experiment is 1/3, then P(failure) is simply 1-1/3 = 2/3."

OpenStudy (anonymous):

so i have divide the number

OpenStudy (mathmale):

Look at the experiment of tossing a fair coin. The coin could land heads up or tails up. there are only those two possible outcomes. Because the coin is a fair one, the probability that you'll get a HEAD is 1/2. The probability that you'll get a TAIL is 1-1/2 = 1/2. In your math problem, there are again only two possible outcomes. If the probability of one outcome is 1/3, then the probability of the other outcome is 1-1/3.